Halloween:

Numerical Problem The pressure in water pipe at the ground floor of a building is 120000 Pa, whereas the pressure on the thurd floor is 30000 Pa. Calculate the height of third floor.

LKylie:

Difference in pressure of water at ground floor and third floor = (120000-30000)=90000 Pa. Let 'h' be the height of third floor. Therefore, Pressure of water due to height h = hqg = h x 1000 x 10 = 10000 h 10000 h = 90000 h = 9m.

Halloween:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @LKylie Difference in pressure of water at ground floor and third floor = (120000-30000)=90000 Pa. Let 'h' be the height of third floor. Therefore, Pressure of water due to height h = hqg = h x 1000 x 10 = 10000 h 10000 h = 90000 h = 9m. \(\color{#0cbb34}{\text{End of Quote}}\) Thnx.